A stock that currently trades for $40 per share is expected to pay a year-end dividend of $2 per share. The dividend is expected to grow at a constant rate over time. The stock has a beta of 1.2, the risk-free rate is 5 percent, and the market risk premium is 5 percent. What is the stock’s expected price seven years from today?
1
Expert's answer
2020-05-11T18:40:16-0400

"k=R(f)+\beta * Risk premium"

k-the rate of yield

R(f)-the risk-free rate

"k=0.05+1.2*0.05=0.11"

"P=DPS*(1+g)\/(k-g)"

P-price of stock

DPS-Dividends Per Share

g-expected growth

"40=2*(1+g)\/(0.11-g)"

"2.2-20g=1+g"

"1.2=21g"

"g=0.057"

P7=2*1.057^7/0.11=26.8